Systems Staff Meeting Notes
August 6, 2007
August 6, 2007
Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Stephen WilliamsAnnouncements
Review Minutes of 7/30/07
- "Bourne Ultimatum" - Tom and Stephen say it is a good movie. Action-packed.
- Tom also saw "Talk to Me" with Don Cheadle - it is very good as well. Tom predicts an Oscar nod for Cheadle.
- Amy survived Vacation Bible School.
- Chris met UF basketball star, Lee Humphrey, in the Gator Shop at the mall this weekend.
- Stephen participated in a 48 Hour Film Project over the weekend. He mostly worked on lighting, although some Systems staff thought 'Sparkles' would have made a great stage name.
- Support Services meeting - they are really scrutinizing the budget. The stat is looking at money that was not spent last year and threatening to take it back. No announcement as to how much the callback would be.
- The conference-call model VOIP phone ordered for the Systems conference room will be switched out with the model ordered for the East conference room.
Ongoing Systems Updates
- Networking item #3 removed.
- Orders - supplies such as batteries and CDs that we redistribute to staff will be special orders placed through Business Services and not come out of our budget.
- Computer inventory - Bobbie has received updates, but every one of them has been incorrect. Mac addresses are wrong, etc. At this rate, it will take a long time for our inventory to be accurate.
- Education has not received their public machines yet.
- Training rescheduled for August 28-31.
- Verify Patron - failover problem. Batch file using Deep Freeze. Amy will test it Wednesday night in Music. This is to fix the folder names for the failover.
Software Projects Status
- Xerox - nothing new.
- Library East VOIP and Gigabit Ethernet upgrade - no word on VOIP. Aug. 11 will be the upgrade to the south tower. All of our servers will be unavailable. The Libraries will all be closed on Saturday. The north tower will be upgraded the following weekend, Aug. 18.
- Education - Media Tech is ready for the electrician to wire the screen.
- Logan is out at VM Ware training.
- DLC - they have a project ongoing that involves a lot of different people and ideas: Logan, Will - now Bill. As of now, we are going to buy them a switch for them to use to network the CopiBook scanners. They will use the 10.227.223 IP range.
- Grover - the Admin tab is almost complete (1/8th of the project). Ticket will be the next section.
- Tracking database - the test version from July 12 is still not finished being tested. Tom is ready to give them a new one, but there is no where to put it.
- Switch to ALEPH v.18 - various problems last week were taken care of: patron files, Oracle database.
Students and Department
- The Food Policy Survey will remain up until fall semester.
- Fundraising site - no word from Barbara Hood.
- Endeca - no browsing; only keyword search. People are bothered by it. Tom has to change all of the links that are pointing to Endeca.
- Issue of dynamic header and footer for front page extension.
Systems Security Management
- Will probably be hiring another new student for the fall semester. His name is Robert, he was recommended by Vera.
- Fall schedule has been worked out with Luiza, Mindy and (hopefully) Robert.
- CNS is asking that we convert all of our 3480 tapes to VTS as they are doing away with the 3480 tape drives. We will have them converted for archival purposes at no charge to us.
- Meeting held.